Find the volume of a cylinder with a diameter of 10 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank)


1
Expert's answer
2020-07-05T18:15:36-0400

From the task condition we find the radius"(r)" and height"(h)":

"r=\frac{d}{2}=5" inches,

"h=3r=15" inches.


The volume "(V)" of the cylinder:

"V=S_{base}h=15\pi r^2=15\pi 5^2=1177.5" inches.


Answer: "V=1177.5" inches.